what is the formula for measuring the price elasticityof supply? suppose the rice of apples



1
Expert's answer
2021-12-01T10:17:03-0500

Formula for price elasticity of supply:

Price Elasticity of supply= percentage change in quantity supplied/percentage change in price.

"=\frac{Percentage Change in Quantity supplied}{Percentage Change in Price}"


"PES=\frac{\frac{1200-1000}{1100}\times100}{\frac{22-20}{21}\times100}"

"PES=1.91=2."

Since price elasticity of supply is greater than 1, the supply is elastic.